# A Comprehensive Survey in LLM(-Agent) Full Stack Safety: Data, Training and Deployment

### v1

@@ Abstract
The remarkable success of Large Language Models (LLMs) has illuminated a promising pathway toward achieving Artificial General Intelligence for both academic and industrial communities, owing to their unprecedented performance across various applications. As LLMs continue to gain prominence in both research and commercial domains, their security and safety implications have become a growing concern, not only for researchers and corporations but also for every nations. Currently, existing surveys on LLM safety primarily focus on specific stages of the LLM lifecycle, e.g., deployment phase or fine-tuning phase, lacking a comprehensive understanding of the entire "lifechain" of LLMs. To address this gap, this paper introduces, for the first time, the concept of "full-stack" safety to systematically consider safety issues throughout the entire process of LLM training, deployment, and eventual commercialization. Compared to the off-the-shelf LLM safety surveys, our work demonstrates several distinctive advantages: (I) Comprehensive Perspective. We define the complete LLM lifecycle as encompassing data preparation, pre-training, post-training (including alignment and fine-tuning, model editing, etc.), deployment and final commercialization. To our knowledge, this represents the first safety survey to encompass the entire lifecycle of LLMs. (II) Extensive Literature Support. Our research is grounded in an exhaustive review of over 800+ papers, ensuring comprehensive coverage and systematic organization of security issues within a more holistic understanding. (III) Unique Insights. Through systematic literature analysis, we have developed reliable roadmaps and perspectives for each chapter. Our work identifies promising research directions, including safety in data generation, alignment techniques, model editing, and LLM-based agent systems. These insights provide valuable guidance for researchers pursuing future work in this field. We provide an up-to-date review of the literature on LLM (agent) safety at https://github.com/bingreeky/full-stack-llm-safety, which can be considered a useful support for both researchers and engineers.
@@ 1 Introduction
Unfortunately, the entire lifecycle of LLMs is constantly confronted with security and safety issues [33, 34, 35, 2025-CodeLM-Security,qu2025prompt]. During the data preparation phase, since LLMs require ample and diverse data, and a significant amount of data is sourced from the Internet and other open-source scenarios, the toxicity in the data and user privacy may seep into the model parameters, triggering crises in the model [36, 37, 38]. The pretraining process of the model, due to its unsupervised nature, unconsciously absorbs these toxic data and privacy information, thereby causing the model’s “genetic makeup” to carry dangerous characteristics and privacy issues [39, 40, 41, 42].
@@
TABLE I: Survey Comparison on LLMs and Agents.
@@
Contribution 1. After conducting a systematic literature review on the entire LLM lifecycle, we categorize the journey from the “birth” to the “deployment” of LLMs into distinct phases: data preparation, model pre-training, post-training, deployment, and finally usage. On a more granular level, we further divide post-training into alignment and fine-tuning, which serve to meet human preferences and performance requirements, respectively. Building upon this, we incorporate model editing and unlearning into our considerations, as methods for efficiently updating the model’s knowledge or parameters, thereby effectively ensuring the model’s usability during deployment. In the deployment phase, we delineate the safety of large models into pure LLM models, which do not incorporate additional modules, and LLM-based agents, which are augmented with tools, memory, and other modules. This framework encompasses the entire cycle of model parameter training, convergence, and solidification.
@@
Figure 1: The overview of the safety of LLM-based agent systems.
@@
Taxonomy. Our article begins with the structural preparation of data. In Section 2, we systematically introduce potential data issues during various model training phases, as well as the currently popular research on data generation. In Section 3, we focus on the security and safety concerns during the pre-training phase, which includes two core modules: data filtering and augmenting. In Section 4, we concentrate on the post-training phase, differing from previous works by incorporating fine-tuning and alignment, which involve attack, defense, and evaluation. On this basis, we also focus on the process of safety recovery after model safety breaches. In Section 5, we observe that models require dynamic updates in real-world scenarios. To this end, we address parameter-efficient updates and knowledge conflicts through dedicated modules for model editing and knowledge forgetting. Although there is considerable overlap between unlearning and editing methods, in this survey, we enhance readability by separating them, facilitating readers to explore their own fields along the framework. Subsequently, in Section 6, we focus on the safety issues after the model parameters are solidified, which share many commonalities with traditional large model security surveys. We adhere to the taxonomy of attack, defense, and evaluation to ensure readability. Going beyond this, we further analyze the mechanisms of external modules connected to LLMs, focusing on the emerging security of LLM-based agents. Finally, in Section 7, we present multiple safety concerns for the commercialization and ethical guidelines, as well as user usage, of LLM-based applications. To provide readers with a comprehensive understanding of our research framework, we dedicate Section 8 to outlining promising future research directions, while Section 9 presents synthesized conclusions and broader implications.
@@ 2.1 Pretraining Data Safety
The pretraining phase of LLMs relies heavily on massive, diverse datasets collected from the Internet [77, 78, 79] or open-source data platforms [80, 81] (e.g., GitHub and Hugging face) providing the foundational “fuel” for their performance. However, this dependence introduces significant safety [82, 83, 84] and privacy risks [85, 86, 87], as the quality, integrity, and safety of the data directly impact the resulting models. This subsection reviews critical threats to pre-training data safety, including data poisoning, privacy leakage, and explores mitigation strategies based on recent literature [88, 81, 86, 89].
@@
Training Data Poisoning. The pre-training phase of LLMs is increasingly recognized as a vulnerable point for data poisoning attacks [39, 40, 90]. These attacks involve the injection of malicious content into training datasets, with the goal of inducing harmful behaviors in the model during inference. Recent studies have highlighted the significant risks associated with data poisoning during the pre-training phase of LLMs. For example, [83] and [84] both highlight that small fractions of poisoned data (as low as 0.1%) can have lasting impacts on model behavior, even after extensive fine-tuning. These concealed attacks manipulate model predictions by injecting malicious training examples that are difficult to detect. Meanwhile, [82] and [91] emphasize the risks of poisoning web-scale datasets, noting that modifying publicly available data (e.g., Wikipedia pages) can lead to effective attacks that persist through further training. The study by Sun et al. [80] shows that code poisoning by simply modifying one variable/function name can enable the code language model for the code search task to make vulnerable code rank in the top 11%.
@@
Privacy leakage. The pre-training phase of language models has become a focal point for discussions on privacy leakage [68, 92, 93, 94, 95, 96]. As these models grow in scale and capability, the risk of inadvertently capturing and leaking personally identifiable information (PII) from their training data becomes more pronounced [41]. [97, 98, 99] have specifically highlighted this concern in the context of LLMs, demonstrating that these models can memorize and reproduce sensitive information through targeted attacks. Data Extraction Attacks such as [100, 101, 102, 103, 104] have shown that even small portions of poisoned data can lead to lasting impacts on model behavior, including the unintentional disclosure of sensitive information. This risk is further underscored by the findings of [39, 40], which emphasize the extent of memorization across different models and the need for robust data management practices to mitigate privacy risks. Meanwhile, Membership Inference Attacks [105, 106, 107, 108], have been shown to be effective in determining whether specific data samples were used during model training in language models, yet recent research [109, 110, 111, 112, 113] indicates that in LLMs, MIA barely outperform random guessing for most settings across varying LLM sizes and domains. Moreover, the research presented in [114, 85] discusses the challenges and applications of protecting data privacy in LLMs, reinforcing the importance of addressing these issues in the development and deployment of these models.
@@ 2.2 Fine-tuning Data Safety
The fine-tuning stage of LLMs has become a critical target for data poisoning attacks, which pose sophisticated threats to model safety [115].
Recent research highlights various vulnerabilities across different fine-tuning approaches, demonstrating how attackers can manipulate training data or inject malicious instructions to compromise model behavior. These risks include:
@@
  • Parameter-Efficient Fine-Tuning Risks. Parameter-efficient fine-tuning (PEFT) techniques [121, 122, 123] also face data poisoning risks.  [124] uncovers stealthy and persistent non-alignment on large language models via backdoor injections. Attackers can subtly alter the model’s alignment by injecting backdoor that remain undetected during the fine-tuning process [125] examines how data poisoning attacks can turn generative models degenerate by introducing poisoned data that degrade the model’s overall performance and leads to the generation of harmful content.
  • Federated Learning Risks. In federated learning, a decentralized training paradigm [126, 127, 128], data poisoning attacks present an even greater challenge due to the distributed nature of the process. Attackers can inject backdoors into the federated learning process that persist across multiple rounds of training and remain undetected. [129] proposes a poisoning attack designed to disrupt the safety alignment of LLMs through fine-tuning a local model on automatically crafted, safety-unaligned data. [130] delves into durable backdoor in federated learning, demonstrating that attackers can create backdoor that are difficult to detect and remove, posing a significant threat to the safety of federated learning models.
@@ 2.3 Alignment Data Safety
  • Human Feedback Stage. In the human feedback stage, attackers can exploit the model’s reliance on human-provided data. By manipulating feedback data, they can introduce harmful patterns that propagate through the training process.
    Recent studies demonstrate three primary attack vectors: (1) [133] develops poisoning techniques using malicious instruction injections that systematically degrade model performance on targeted tasks. (2) [134, 135] engineer universal jailbreak backdoor through feedback manipulation, creating persistent vulnerabilities that bypass safety constraints when triggered by specific prompts. (3) [136] crafts deceptive feedback that induces incorrect or harmful outputs.
  • Reinforcement Learning from Human Feedback (RLHF) Stage.
    In the RLHF stage, the integrity of the model’s learning process can be compromised through the poisoning of reward models  [1, 137, 138, 139, 140, 141]. A critical example is the RankPoison attack introduced by [142], which manipulates reward signals by strategically corrupting human preference datasets. Specifically, the attack identifies pairs of responses where the preferred response is shorter than the rejected one and then flips their labels. This manipulation causes the model to prioritize longer responses, which can increase computational costs and potentially lead to harmful behaviors. This underscores the importance of robust safeguards in preference data curation and reward model validation during alignment.
@@ 2.4 Safety in Data Generation
Data Generation in the Lifecycle of LLMs. Data synthesis has become an indispensable component of every phase in the LLM ecosystem: in the (i) pre-training stage, LLM-based data generation is often referred to as model distillation, where corpora generated by larger models serve as training data for smaller models, as seen in Phi-1 [143], Phi-1.5 [144], and AnyGPT [145], among others. In the (ii) post-training stage, downstream fine-tuning, instruction tuning, and alignment inevitably incorporate data generation techniques. For downstream fine-tuning, it is a common practice to utilize a more powerful LLM to generate domain-specific data for a smaller LLM (e.g., Chinese medical knowledge in [146], multiple-choice question answering in [147], mathematical reasoning in [148], and clinical text data [149]) to enhance its domain-specific capabilities. It is also empirically validated that LLM-generated data (e.g., action trajectories, question-answer pairs) can be beneficial for improving the reasoning [150], planning, function calling [151] abilities. For instruction tuning, some approaches employ powerful LLMs to generate instruction-tuning data, such as Evol-Instruct from WizardLM [152] and Orca [153], while others adopt self-instruct techniques like Self-Instruct [154] and Self-Translate [155]. For alignment, models such as Beavertails [156], PRM800K [157], and WebGPT [158] extensively rely on LLMs for question/response generation, preference ranking for preference dataset synthesis.
@@
Safety Issues and Mitigation. Despite its success, data generation inevitably introduces additional uncertainties and security risks throughout the LLM lifecycle, primarily in the following aspects: (1) Privacy, where synthetic data generation poses risks of amplifying privacy leakage due to the memorization of sensitive training samples and inadequate anonymization [159], particularly in privacy-sensitive applications such as medical text processing [160] and disease diagnosis [161]. (2) Bias and Fairness, as LLMs inherently exhibit societal biases [162] (e.g., gender stereotypes in job descriptions), and the data they generate may further exacerbate these biases [163, 164]. This issue can and should be mitigated during the data filtering process using existing LLM debiasing techniques [165, 166, 167]. (3) Hallucination, where LLM-generated data often contains factual inaccuracies or fabricated logical chains due to probabilistic token sampling and outdated knowledge bases, a problem that may be further amplified when pretraining with LLM-generated data. Potential solutions include filtering generated data using existing hallucination detection techniques [168, 169]. (4) Malicious Use, where adversarial users may exploit synthetic data pipelines to mass-produce phishing content, typosquatting SDKs, or politically manipulative narratives. (5) Misalignment, where RLHF in LLM training can be compromised by selectively manipulating data samples in the preference dataset [170].

### v2

@@ Abstract
The remarkable success of Large Language Models (LLMs) has illuminated a promising pathway toward achieving Artificial General Intelligence for both academic and industrial communities, owing to their unprecedented performance across various applications. As LLMs continue to gain prominence in both research and commercial domains, their security and safety implications have become a growing concern, not only for researchers and corporations but also for all nations. Currently, existing surveys on LLM safety primarily focus on specific stages of the LLM lifecycle, e.g., deployment phase or fine-tuning phase, lacking a comprehensive understanding of the entire “lifechain” of LLMs. To address this gap, this paper introduces, for the first time, the concept of “full-stack” safety to systematically consider safety issues throughout the entire process of data, training (pre-training, post-training), deployment (deployment and final commercialization). Compared to the off-the-shelf LLM safety surveys, our work demonstrates several distinctive advantages: (I) Comprehensive Perspective. We define the complete LLM lifecycle as encompassing data preparation, pre-training, post-training (including alignment and fine-tuning, model editing, etc.), deployment and final commercialization. To our knowledge, this represents the first safety survey to encompass the entire lifecycle of LLMs. (II) Extensive Literature Support. Our research is grounded in an exhaustive review of over 900+ papers, ensuring comprehensive coverage and systematic organization of safety issues within a more holistic understanding. (III) Unique Insights. Through systematic literature analysis, we develop reliable roadmaps and perspectives for each chapter. Our work identifies promising research directions, including safety in data generation, alignment techniques, model editing, and LLM-based agent systems. These insights provide valuable guidance for researchers pursuing future work in this field. We provide an up-to-date review of the literature on LLM (agent) safety at https://github.com/bingreeky/full-stack-llm-safety, which can be considered a useful support for both researchers and engineers.
@@ 1 Introduction
Unfortunately, the entire lifecycle of LLMs is constantly confronted with security and safety issues [33, 34, 35, 36, 37]. During the data preparation phase, since LLMs require ample and diverse data, and a significant amount of data is sourced from the Internet and other open-source scenarios, the toxicity in the data and user privacy may seep into the model parameters, triggering crises in the model [38, 39, 40]. The pretraining process of the model, due to its unsupervised nature, unconsciously absorbs these toxic data and privacy information, thereby causing the model’s “genetic makeup” to carry dangerous characteristics and privacy issues [41, 42, 43, 44].
@@
TABLE I: Survey Comparison on LLMs and Agents settings.
@@
Contribution 1. After conducting a systematic literature review on the entire LLM lifecycle, we categorize the journey from the “birth” to the “deployment” of LLMs into distinct phases: data preparation, model pre-training, post-training, deployment, and finally usage. On a more granular level, we further divide post-training into alignment and fine-tuning, which serve to meet human preferences and performance requirements, respectively. Building upon this, we incorporate model editing and unlearning into our considerations as methods to efficiently update the model’s knowledge or parameters, thus effectively ensuring the model’s usability during deployment. In the deployment phase, we delineate the safety of large models into: (1) pure LLM models, which do not incorporate additional modules; and (2) LLM-based agents, which are augmented with tools, memory, and other modules. This framework encompasses the entire cycle of model parameter training, convergence, and solidification.
@@
Figure 1: We present a systematic taxonomy while enumerating notable works (2022-2025) and their institutional affiliations.
@@
Taxonomy. Our article begins with the structural preparation of data. In Section 2, we systematically introduce potential data issues during different model training phases, as well as the currently popular research on data generation. In Section 3, we focus on the security and safety concerns during the pre-training phase, which includes two core modules: data filtering and augmenting. In Section 4, we concentrate on the post-training phase, differing from previous works by incorporating fine-tuning and alignment, which involve attack, defense, and evaluation. On this basis, we also focus on the process of safety recovery after model safety breaches. In Section 5, we observe that models require dynamic updates in real-world scenarios. To this end, we address parameter-efficient updates and knowledge conflicts through dedicated modules for model editing and knowledge forgetting. Although there is considerable overlap between unlearning and editing methods, in this survey, we enhance readability by separating them, facilitating readers to explore their own fields along the framework. Subsequently, in Section 6, we focus on the safety issues after the model parameters are solidified, which share many commonalities with traditional large model security surveys. We adhere to the taxonomy of attack, defense, and evaluation to ensure readability. Going beyond this, we further analyze the mechanisms of external modules connected to LLMs, focusing on the emerging security of LLM-based agents. Finally, in Section 7, we present multiple safety concerns for the commercialization and ethical guidelines, as well as user usage, of LLM-based applications. To provide readers with a comprehensive understanding of our research framework, we dedicate Section 8 to outlining promising future research directions, while Section 9 presents synthesized conclusions and broader implications.
@@ 2.1 Pretraining Data Safety
The pretraining phase of LLMs relies heavily on massive, diverse datasets collected from the Internet [78, 79, 80] or open-source data platforms [81, 82] (e.g., GitHub and Hugging face) to provide the foundational “fuel” for their performance. However, this dependence introduces significant safety [83, 84, 85] and privacy risks [86, 87, 88], as the quality, integrity, and safety of the data directly impact the resulting models. This subsection reviews critical threats to pre-training data safety, including data poisoning, privacy leakage, and explores mitigation strategies based on recent literature [89, 82, 87, 90].
@@
Training Data Poisoning. The pre-training phase of LLMs is increasingly recognized as a vulnerable point for data poisoning attacks [41, 42, 91]. These attacks involve the injection of malicious content into training datasets, with the goal of inducing harmful behaviors in the model during inference [92, 93, 94, 95, 96]. Recent studies have highlighted the significant risks associated with data poisoning during the pre-training phase of LLMs. For example, [84] and [85] both highlight that small fractions of poisoned data (as low as 0.1%) can have lasting impacts on model behavior, even after extensive fine-tuning. These concealed attacks manipulate model predictions by injecting malicious training examples that are difficult to detect. Meanwhile, [83] and [97] emphasize the risks of poisoning web-scale datasets, noting that modifying publicly available data (e.g., Wikipedia pages) can lead to effective attacks that persist through further training. The study by Sun et al. [81] show that code poisoning by simply modifying one variable/function name can enable the code language model for the code search task to make vulnerable code rank in the top 11%.
@@
Privacy leakage. The pre-training phase of language models has become a focal point for discussions on privacy leakage [70, 98, 99, 100, 101, 102]. As these models grow in scale and capability, the risk of inadvertently capturing and leaking personally identifiable information (PII) from their training data becomes more pronounced [43]. [103, 104, 105] have specifically highlighted this concern in the context of LLMs, demonstrating that these models can memorize and reproduce sensitive information through targeted attacks. Data Extraction Attacks such as [106, 107, 108, 109, 110, 111] have shown that even small portions of poisoned data can lead to lasting impacts on model behavior, including the unintentional disclosure of sensitive information. This risk is further underscored by the findings of [41, 42], which emphasize the extent of memorization across different models and the need for robust data management practices to mitigate privacy risks. Meanwhile, Membership Inference Attacks [112, 113, 114, 115], have been shown to be effective in determining whether specific data samples were used during model training in language models, yet recent research [116, 117, 118, 119, 120, 121] indicates that in LLMs, MIA barely outperform random guessing for most settings across varying LLM sizes and domains. Moreover, the research presented in [122, 86] discusses the challenges and applications of protecting data privacy in LLMs, reinforcing the importance of addressing these issues in the development and deployment of these models.
Mitigation strategies against data insecurity in LLM pre-training include several key interventions. To address toxic content, custom classifiers trained on safety datasets are employed to detect and filter pre-training data [123, 89, 124]. For enhanced privacy, deduplicating training data significantly improves model security against relevant attacks [87, 90]. Furthermore, safety awareness is cultivated during pre-training by managing model outputs through safety plans or by marking and removing unsafe generations [125, 123, 126, 82], leading to safer and more executable planning capabilities.
Mitigation measures. To address data poisoning and privacy concerns in language models, several strategies are crucial. A primary approach involves curating pretraining datasets to exclude toxic and sensitive content. [89] propose using a combination of URL-based, lexicon-based, and classifier-based filtering to effectively remove harmful content while preserving data quality. Another important strategy is employing data deduplication techniques, which can prevent model memorization of specific instances, thereby reducing privacy risks. [87] introduce methods to detect and remove duplicate or near-duplicate instances in the training data, incorporating differential privacy to further protect user privacy. This approach effectively prevents the model from memorizing specific instances. In addition, developing robust defenses against data poisoning is vital to ensure that models are less susceptible to manipulation through malicious data injection. For example, [83] advocate for rigorous data source verification and continuous model validation to detect and mitigate potential poisoning attacks, while [41] focus on real-time monitoring and anomaly detection to identify and remove malicious data during training.
@@ 2.2 Fine-tuning Data Safety
Data safety in the fine-tuning stage has emerged as a critical concern in the development of LLMs, with data poisoning attacks presenting particularly sophisticated threats to LLMs [127]. Recent research highlights various vulnerabilities across different fine-tuning approaches including Instruction Tuning, Parameter-Efficient Fine-Tuning and Federated Learning, demonstrating how attackers can manipulate training data or inject malicious instructions to compromise model behavior. These risks include:
@@
  • Parameter-Efficient Fine-Tuning Risks. Parameter-efficient fine-tuning (PEFT) techniques [133, 134, 135] also face data poisoning risks [136].  [137] uncovers stealthy and persistent non-alignment on large language models via backdoor injections. Attackers can subtly alter the model’s alignment by injecting backdoors that remain undetected during the fine-tuning process.  [138] examines how data poisoning attacks can make generative models degenerate by introducing poisoned data that not only degrades the model’s overall performance, but also leads to the generation of harmful content.
  • Federated Learning Risks. Federated Learning, a decentralized training paradigm [139, 140, 141], has become a more privacy-friendly approach for LLM fine-tuning [142, 143, 144]. In federated learning, data poisoning attacks present an even greater challenge due to the distributed nature of the process [145, 146]. Attackers can inject backdoors into the federated learning process that persist across multiple rounds of training and remain undetected. [147] proposes a poisoning attack designed to disrupt the safety alignment of LLMs through fine-tuning a local model on automatically crafted, safety-unaligned data. [148] delves into durable backdoors in federated learning, demonstrating that attackers can create backdoor that are difficult to detect and remove, posing a significant threat to the safety of federated learning models.
@@ 2.3 Alignment Data Safety
  • Human Feedback Stage. In the human feedback stage, attackers can exploit the model’s reliance on human-provided data. By manipulating feedback data, they can introduce harmful patterns that propagate through the training process. Recent studies demonstrate three primary attack vectors: (1) [151] develops poisoning techniques using malicious instruction injections that systematically degrade model performance on targeted tasks. (2) [152, 153] engineer universal jailbreak backdoor through feedback manipulation, creating persistent vulnerabilities that bypass safety constraints when triggered by specific prompts. (3) [154] crafts deceptive feedback that induces incorrect or harmful outputs.
  • Reinforcement Learning from Human Feedback (RLHF) Stage. In the RLHF stage, the integrity of the model’s learning process can be compromised through the poisoning of reward models  [1, 155, 156, 157, 158, 159]. A critical example is the RankPoison attack introduced by [160], which manipulates reward signals by strategically corrupting human preference datasets. Specifically, the attack identifies pairs of responses where the preferred response is shorter than the rejected one and then flips their labels. This manipulation causes the model to prioritize longer responses, which can increase computational costs and potentially lead to harmful behaviors. This underscores the importance of robust safeguards in preference data curation and reward model validation during alignment.
@@ 2.4 Safety in Data Generation
Data Generation in the Lifecycle of LLMs. Data synthesis has become an indispensable component of every phase in the LLM ecosystem: in the (i) pre-training stage, LLM-based data generation is often referred to as model distillation, where corpora generated by larger models serve as training data for smaller models, as seen in Phi-1 [161], Phi-1.5 [162], and AnyGPT [163], among others. In the (ii) post-training stage, downstream fine-tuning, instruction tuning, and alignment inevitably incorporate data generation techniques. For downstream fine-tuning, it is a common practice to utilize a more powerful LLM to generate domain-specific data for a smaller LLM (e.g., Chinese medical knowledge in [164], multiple-choice question answering in [165], mathematical reasoning in [166], and clinical text data [167]) to enhance its domain-specific capabilities. It is also empirically validated that LLM-generated data (e.g., action trajectories, question-answer pairs) can be beneficial for improving the reasoning [168, 169], planning, function calling [170] abilities. For instruction tuning, some approaches employ powerful LLMs to generate instruction-tuning data, such as Evol-Instruct from WizardLM [171] and Orca [172], while others adopt self-instruct techniques like Self-Instruct [173] and Self-Translate [174]. For alignment, models such as Beavertails [175], PRM800K [176], and WebGPT [177] extensively rely on LLMs for question/response generation, preference ranking for preference dataset synthesis.
@@
Safety Issues and Mitigation. Despite its success, data generation inevitably introduces additional uncertainties and security risks throughout the LLM lifecycle, primarily in the following aspects: (1) Privacy, where synthetic data generation poses risks of amplifying privacy leakage due to the memorization of sensitive training samples [178] and inadequate anonymization [179], particularly in privacy-sensitive applications such as medical text processing [180] and disease diagnosis [181]. (2) Bias and Fairness, as LLMs inherently exhibit societal biases [182] (e.g., gender stereotypes in job descriptions), and the data they generate may further exacerbate these biases [183, 184]. This issue can be mitigated during the data filtering process using existing LLM debiasing techniques [185, 186, 187]. (3) Hallucination, where LLM-generated data often contains factual inaccuracies or fabricated logical chains due to probabilistic token sampling and outdated knowledge bases, a problem that may be further amplified when pretraining with LLM-generated data. Potential solutions include filtering generated data using existing hallucination detection techniques [188, 189]. (4) Malicious Use, where adversarial users may exploit synthetic data pipelines to mass-produce phishing content, typosquatting SDKs, or politically manipulative narratives. (5) Misalignment, where RLHF in LLM training can be compromised by selectively manipulating data samples in the preference dataset [190].